# SELFI: Autonomous Self-Improvement with Reinforcement Learning for Social Navigation

### v1

@@ IV-A Offline Learning with SACSoN
We briefly describe the objective and the learning procedure in SACSoN. Details are shown in the original paper [18].
Offline model-based objectives: As shown in Fig. 3, we represent the control policy by an encoder $g_{\phi}$ coupled to an actor network $\pi_{\theta}$. We train the policy to maximize the model-based objective $J_{\text{sacson}}$ [18] given by four terms:
$J_{\text{sacson}}(s,\tau):=\sum_{t=0}^{H-1}\hat{r}^{\tt{pose}}_{t}+\hat{r}^{\tt{geom}}_{t}+\hat{r}^{\tt{ped}}_{t}+\hat{r}^{\tt{reg}}_{t}$ (9)
@@
where the $\hat{r}^{\tt{pose}}$ reward encourages goal-reaching behavior, $\hat{r}^{\tt{geom}}$ [17] penalizes collision with static obstacles (via signed point-cloud distance), $\hat{r}^{\tt{ped}}$ [18] is to learn socially unobtrusive behavior and $\hat{r}^{\tt{reg}}$ acts as a regularization term to encourage smooth motion. All objectives in Eqn 9 are differentiable (through the forward model) with respect to the action sequence $\tau$ and can therefore calculate the gradient of the encoder $g_{\phi}$ and the actor network $\pi_{\theta}$ via $\tau$ and learn them.
@@
IV-B Online Learning with SELFI
We demonstrate an instantiation of SELFI for the socially-compliant navigation task following the design in Sec. III-B and Sec. III-C. Here, we describe the specific procedure we use for online RL finetuning.
Learning setting: To obtain the target robotic behavior as fast as possible during online learning, SELFI fine-tunes the actor $\pi_{\theta}$ with a frozen encoder $g_{\phi}$ as shown in Fig. 3. It corresponds to defining the extracted feature from the encoder as $s$ in formulation of SELFI. To estimate $\bar{Q}$, we feed the feature $s$ and the sequence of action $\tau$ to the critic, as shown in bottom part of Fig. 3. By sharing the encoder with the actor, we efficiently learn both actor and critic. In the TD error calculation, we assume $J(s,\tau)\approx\gamma J(s^{\prime},\tau^{\prime})$ to train the small critic network during the brief online learning phase.
Reward design: The reward $r$ is designed to incentivize smoothness and progress towards the goal, while avoiding collisions:
$r(s,a)=\vec{v}\cdot\hat{g}+C_{s}+C_{d},$ (10)
The first term maximizes velocity towards the next goal, where $\hat{g}:=[x_{g},y_{g},\theta_{g}]$ and $\vec{v}:=[v,0,\omega]$ are goal direction and velocity vectors expressed in the robot’s current frame [46]. To obtain the local goal pose $g$, we build an approximate localization system that incorporates visual odometry and AR markers along the robot’s trajectory. For more details, please refer to the appendix. $v$ and $\omega$ are the linear and the angular velocity commands for the two-wheel-drive robot in $\hat{a}_{0}$. We use $C_{s}$ and $C_{d}$ to denote the rewards for avoiding static obstacles (environment geometry) and dynamic obstacles (pedestrians) respectively:
@@ IV-B Online Learning with SELFI
$C_{d}$ $=$ $\begin{cases}-C_{h}\hskip 14.22636pt(\mbox{if}\hskip 4.2679ptd_{h}<0.5)\\ \hskip 8.53581pt0.0\hskip 14.22636pt(\mbox{otherwise})\\ \end{cases}.$ (12)
@@
The robot triggers a collision event using the robot’s bumper sensor, and a bumpy event (caused by an uneven floor) when the measured acceleration exceeds a fixed threshold. To detect personal space violations, we estimate the distance $d_{h}$ to the closest pedestrians using a combination of semantic segmentation [38, 1] and monocular depth estimation [14].
Others: We set the discount factor $\gamma=0.97$, accounting for long trajectories of human-robot interaction behaviors. In addition, the workstation sends the policy model parameters $\theta$ to the robot every 50 training steps (approx. 1 minute wall clock time). All other parameters follow the authors’ implementation of TD3 [10] and SACSoN [18].
IV-C Robotic system
Hardware setup: Figure 4 shows the overview of our prototype robot. We use an omnidirectional camera to observe $\{I_{t}\}_{t=-M\ldots 0}$ and $I^{g}$. This allows us to observe a ${}^{\circ}$ view for capturing the pedestrians even behind the robot. The robot is equipped with an NVIDIA Jetson Orin AGX onboard computer, which runs inference of trained models at 3 Hz. We use two additional cameras to estimate visual odometry and to detect long-term localization fiducials, following the setup of Hirose et al. [18]. We use an IMU to measure bumpiness and uneven terrain and a bumper sensor to detect collisions.
In addition to on-robot compute, we use a workstation for fast, online training. The workstation is equipped with an Intel i9 CPU, 96GB RAM, and an NVIDIA RTX 3090ti GPU.
Figure 4: Overview of the prototype robot [33]. We use the Ricoh Theta S omnidirectional camera and run inference on a Nvidia Orin AGX. Yellow description boxes are components used in online learning only.
Figure 5: Three environments on online training and evaluation. We conduct online training in three different challenging environments, [a] Environment 1 is the open space facing restrooms, elevator hall and café space, [b] Environment 2 is in the entrance hall with many pedestrians, and, [c] Environment 3 is along the office area with narrow corridors. Environment 1 and 3 have many glass walls, which are difficult for collision avoidance and cause lighting condition changes.
Figure 6: Small obstacles in online training and evaluation. We randomly place obstacles during training (left) and conduct the evaluation with unseen obstacles (right).
V Evaluation
@@
    Does SELFI reduce interventions during the fine-tuning process (without performance degradation)?
  3. Q3.
    Can SELFI be improved by fine-tuning on data from the target environment during pre-training?
V-A Evaluation setup
@@
Residual RL [19]: In residual RL, the policy is given by the sum of a base policy and a learned policy, $a=\pi^{\text{base}}(s)+\pi^{\text{RL}}(s)$, where $\pi^{\text{base}}$ is the pre-trained control policy and $\pi^{\text{RL}}$ is the actor trained with online RL (TD3). We evaluate two choices for $\pi^{\text{base}}$: (1) the pre-trained control policy maximizing only $\sum_{t=0}^{H-1}\hat{r}^{\tt{pose}}_{t}$ to simply move towards the goal position, and (2) the pre-trained SACSoN policy, maximizing the total objective $J_{\text{sacson}}$. We label the latter as Residual ${}^{\dagger}$. Residual RL provides an alternative way to combine prior policies with online model-free RL, and therefore represents a natural prior method for comparing with SELFI.
@@
All baselines use the same network structure, except that single-step methods (all except Ours and SACSoN with fine-tuning) predict only a single action $a$ rather than a sequence $\tau$. Unless specified, all RL-based methods use the same reward.
@@
In these three environments, we design the looped trajectories, which the last node is same pose as the initial node, as shown by red lines in Fig. 5, and feed the first goal image when arriving at the last node to continuously train the control policy online. During online training, we randomly place small objects shown in Fig. 6[a] to learn the collision avoidance behavior for the small objects. Since Environment 1 includes many challenges for our task, we evaluate all methods in this environment and conduct the overall evaluations with the selected baselines in three environments.
V-B Performance analysis of the fine-tuned control policy
@@
We compare all of the methods, but first we compare to the two strongest baselines, Residual RL and SACSoN with fine-tuning (as selected according to the experiments in Section V-D). We conduct online learning in each environment and stop training when the robot run 15 laps or two hours. It is difficult to conduct extensive and reproducible comparisons in the highly populated natural environments, because they are uncontrolled, with pedestrians walking in and out of the scene at random. Therefore, to conduct a more controlled and reproducible comparison with each of the baselines, addressing Q1, we focus on the “organized” environments specifically.

### v2

@@ 4.1 Offline Learning with SACSoN
We briefly describe the learning procedure in SACSoN. Details are shown in the original paper [5].
@@
Offline model-based objectives: As shown in Fig. 2, we represent the control policy by an encoder $g_{\phi}$ coupled to an actor network $\pi_{\theta}$. We train the policy to maximize the model-based objective [5], $J_{\text{sacson}}(s,\tau):=\sum_{t=0}^{H-1}\hat{r}^{\tt{pose}}_{t}+\hat{r}^{\tt{geom}}_{t}+\hat{r}^{\tt{ped}}_{t}+\hat{r}^{\tt{reg}}_{t}$ where the $\hat{r}^{\tt{pose}}$ reward encourages goal-reaching behavior, $\hat{r}^{\tt{geom}}$ [25] penalizes collision with static obstacles (via signed point-cloud distance), $\hat{r}^{\tt{ped}}$ [5] is to learn socially unobtrusive behavior and $\hat{r}^{\tt{reg}}$ acts as a regularization term to encourage smooth motion. All objectives in $J_{\text{sacson}}$ are differentiable with respect to the action sequence $\tau$ and can therefore calculate the gradient of $g_{\phi}$ and $\pi_{\theta}$ via $\tau$ and learn them.
@@
4.2 Online Learning with SELFI
We demonstrate an instantiation of SELFI for the socially-compliant navigation, following Sec. 3.2 and Sec. 3.3. Here, we describe the specific procedure we use for online RL finetuning.
Learning setting: To obtain the target robotic behavior as fast as possible during online learning, SELFI fine-tunes the actor $\pi_{\theta}$ with a frozen encoder $g_{\phi}$ as shown in Fig. 2. It corresponds to defining the extracted feature from the encoder as $s$ in formulation of SELFI. To estimate $\bar{Q}$, we feed the feature $s$ and the sequence of action $\tau$ to the critic, as shown in right part of Fig. 2. By sharing the encoder with the actor, we efficiently learn both actor and critic. In the TD error calculation, we assume $J(s,\tau)\approx\gamma J(s^{\prime},\tau^{\prime})$ to train the small critic network during the brief online learning phase.
Reward design: The reward is designed as $r(s,a)=\vec{v}\cdot\hat{g}+C_{s}+C_{d}$ to incentivize smoothness and progress towards the goal, while avoiding collisions: The first term maximizes velocity towards the next goal, where $\hat{g}:=[x_{g},y_{g},\theta_{g}]$ and $\vec{v}:=[v,0,\omega]$ are goal direction and velocity vectors expressed in the robot’s current frame [13]. To obtain the local goal pose $g$, we build an approximate localization system that incorporates visual odometry and AR markers along the robot’s trajectory. The specific choice of AR markers is a design decision to facilitate easy online learning, and other mechanisms for localization based on visual odometry(see Appendix F). $v$ and $\omega$ are the linear and the angular velocity commands for the two-wheel-drive robot in $\hat{a}_{0}$. We use $C_{s}$ and $C_{d}$ to denote the rewards for avoiding static obstacles and dynamic obstacles (pedestrians) respectively:
@@ 4.2 Online Learning with SELFI
The robot triggers a collision event using the robot’s bumper sensor, and a bumpy event (caused by an uneven floor) when the measured acceleration exceeds a fixed threshold. Note that these sensors are not mandatory and can be substituted by the other sensors commonly used in navigation [13]. To detect intimate distance violations, we estimate the distance $d_{h}$ to the closest pedestrians using a combination of semantic segmentation [67, 68] and monocular depth estimation [69]. $r_{r}=$ 0.5 m is the radius of the circular robot footprint with margin.
Others: We set the discount factor $\gamma=0.97$, accounting for long trajectories of human-robot interaction behaviors. In addition, the workstation sends the policy model parameters $\theta$ to the robot every 50 training steps (approx. 1 minute wall clock time). Since we conduct online learning for maximum of two hours, the maximum number of training steps is about 6000. On online learning, we set the batch size as 76. The learning rate of Adam optimizer is set as 0.0001. All other parameters follow the authors’ implementation of TD3 [18] and SACSoN [5].
Table 1: Closed-loop Evaluation of trained control policies. IDV is intimate distance violation duration, NCO is near-collision duration, and UFS is duration on uneven floor surface, CP is the number of collision for pedestrians, CO is the number of collision for the tiny objects, Int is the number of interventions by teleoperators, SPL is Success weighted by Path Length [70] and STL is Success weighted by Time Length [71]. SPL and STL are calculated by assuming that the robot reaches the goal position even regardless of there being a human intervention. * indicates using the ground truth goal pose for generating the velocity commands.
Method IDV [s] $\downarrow$ NCO [s] $\downarrow$ UFS [s] $\downarrow$ CP $\downarrow$ [#] CO $\downarrow$ [#] Int $\downarrow$ [#] SPL $\uparrow$ STL $\uparrow$
Sampling-based motion planning* 21.512 11.633 10.212 1.333 3.333 0.733 0.808 0.652
Residual RL 28.731 12.365 8.073 3.200 4.333 2.067 0.817 0.641
SACSoN with fine-tuning 18.018 11.229 3.578 0.800 2.667 0.733 0.838 0.596
Ours 7.978 3.938 3.067 0.200 1.400 0.267 0.918 0.739
5 Evaluation
@@
    Does SELFI reduce interventions during the fine-tuning process (without degradation)?
Please see Appendix A for more details on a prototype robot and whole navigation system for our evaluation. And we show the details of our online training and evaluation setup in Appendix B.
Performance analysis of the fine-tuned control policy: We compare to the strongest baselines, Residual RL and SACSoN with fine-tuning. We fine-tune our method as well as these baselines and evaluate the fine-tuned policy. In addition, we evaluate Sampling-based motion planning as the classical motion planning. The details of these baseline methods and the online training setups are shown in Appendix B and G. In evaluation, it is difficult to conduct extensive and reproducible comparisons in the highly populated natural environments, because they are uncontrolled, with pedestrians walking in and out of the scene at random. Therefore, to conduct a more controlled and reproducible comparison with each of the baselines, addressing Q1, we focus on the “organized” environments specifically.
We run our robot five laps each in three different environments, with a control policy fine-tuned by our methods and two selected baselines. For more details regarding the experimental setups, please refer to the appendix C.1. Table 1 shows quantitative analysis with our method and selected the three strongest baselines (The measured metrics are shown on the caption).
[…]
To evaluate how well our method behaves around humans, we also conduct additional experiments with twelve human subjects. We recruited twelve subjects from among graduate students, visiting scholars, and staff members on campus. We consider the balance to be as diverse in gender (6 male, 6 female), professional-level (7 student, 5 non-student), and origin (4 North America, 5 Asia, 2 Europe, 1 Latin America) as possible. For more details, please refer to the appendix C.2. Following [72], subjects were asked nine qualitative questionnaires [72] with a 7-point Likert scale from “Not at all” to “Very much” after each experiment. To mitigate bias in favor of the first method, we present the questionnaires [72] in Fig. 3 before running evaluations with the first method.
Figure 3: Evaluation of socialness by human rating. The height of each bar indicates the mean and the range line indicates the standard deviation. Larger is better in all ratings. * indicates the statistical significance of our method on the t-test with $p<$ 0.05. $\dagger$ indicates scale reversed for analyses.
Figure 4: Visualization of the robot behavior when interacting with the pedestrian. Robots and pedestrians with the same color indicate the same time. The black dotted lines indicate the time of closest proximity.
Figure 3 shows the means and standard deviations of the scores from this study. Note that we flip the score for the last question such that higher is better for the entire bar plot. Residual RL often violates the intimate distance and occasionally collides with the human subject. As the result, Residual RL has the worst score. SACSoN with fine-tuning shows similar behavior in navigation and the closest score to our method. However, our method performs better overall. The statistical significance of our method is confirmed in four questions, (4), (6), (8) and (9) on the t-test with $p<$ 0.05.
Figure 4 visualizes the robot behavior when the pedestrian and robot pass each other. In these time lapse illustrations, the color of the pedestrian and robot indicates the timestep (i.e., a yellow pedestrian and a yellow robot indicate the same point in time). In addition, we use a black dotted line for the robot and the pedestrians when robots and people are in closest proximity. In all cases, the pedestrian gets stuck in front of or behind the robot running the baseline method (right), and the robot penetrates the intimate distance for an extended length of time. Especially in case A and D, the baseline fails to reach the goal position and collides with obstacles. With our method, an evasive maneuver is initiated at an early stage and succeeds in smoothly passing a pedestrian without getting stuck (left). Although the robot is close to the pedestrians when passing, our control policy minimizes how long the robot penetrates the pedestrian’s intimate distance.
Figure 5: Visualization of the robot behavior of avoidance for small obstacles and uneven floor mat.
Figure 5 shows the time lapsed images when avoiding the small unseen objects and the uneven rubber mat. Our method naturally avoids colliding with the small objects, though this presents a challenge for the initial SACSoN policy. In addition, our methods avoids traveling on the uneven mat since we give a negative reward $C_{s}$ during online training. Please see our supplemental materials for more details.
Interventions during online training: Reducing human interventions is important in online training to enable autonomous adaptation in the real world. However, it is known that the data distribution shift between offline and online training causes performance degradation, leading navigation failures and a lot of interventions to keep online learning in navigation. To evaluate the online training process, we count the number of human interventions in each navigation loop for Q2.
Figure 6: The number of intervention on online learning in three different environments. The lines indicate the mean and the areas indicate the standard deviation. The horizontal axis indicates the number of times the robot laps the loop reference.
Please see additional evaluation with more baselines in Appendix E and supplemental videos.
6 Discussion
We proposed an online self-improving method, SELFI, to quickly fine-tune a control policy pre-trained with model-based learning. SELFI combines model-based learning and model-free RL its training objectives to take advantage of the best parts of both approaches. The same objectives used in offline learning are introduced into online learning to stabilize the learning process. The performance of the pre-trained policies are improved via Q-functions from online model-free RL.
In the evaluation, SELFI was implemented to fine-tune the SACSoN policy [5] for vision-based navigation. SELFI enables us to quickly learn complex robotic behavior, such as pre-emptive collision avoidance for pedestrians, collision avoidance for the small and transparent obstacles, and preferences for traversing on smooth surfaces. These behaviors are difficult to learn on offline training due to the modeling errors and data distribution shift. In addition, compared to various baseline methods, SELFI did not require much human intervention during online learning. The performance of the trained control policy by SELFI is also visualized in the supplemental materials.
While our method enables us to quickly fine-tune a pre-trained policy, it has some limitations. For effective online learning, the balance between objectives from model-based learning and the learned Q-function is important, but this balance cannot be predicted in advance and requires some trial-and-error with real robots. Although the reward for socialness is given only for intimate distance violations, human-in-the-loop online learning with human evaluations can lead to better behavior. And a more diverse multiple public datasets on offline training is practically required for robustness.
Acknowledgments
We consulted the Committee for Protection of Human Subjects at our home institute, and it was determined that the study does not meet the definition of research with human subjects as outlined in Federal Regulations at 45 CFR 46.102. This research was supported by Berkeley DeepDrive at the University of California, Berkeley, and Toyota Motor North America. And, this work was partially supported by ARL DCIST CRA W911NF-17-2-0181 and ARO W911NF-21-1-0097. We thank Qiyang Li and Mitsuhiko Nakamoto for advising the mathematical formulation. We would like to express their gratitude to Roxana Infante, Ami Katagiri, Pranav Atreya, Stephan Allenspach, Lydia Ignatova, Toru Lin, Charles Xu, Kosuke Tahara, Katsuhiro Kutsuki and Catherine Glossop for their valuable assistance in evaluating SELFI.
Appendix
Appendix A Robotic system
Hardware setup: Figure 7[a] shows the overview of our prototype robot. We use an omnidirectional camera to observe $\{I_{t}\}_{t=-M\ldots 0}$ and $I^{g}$. This allows us to observe a 360 view for capturing the pedestrians even behind the robot. The robot is equipped with an NVIDIA Jetson Orin AGX onboard computer, which runs inference of trained models at 3 Hz. We use two additional cameras to estimate visual odometry and to detect long-term localization fiducials, following the setup of Hirose et al. [5]. We use an IMU to measure bumpiness and uneven terrain and a bumper sensor to detect collisions. In addition to on-robot compute, we use a workstation for fast, online training. The workstation is equipped with an Intel i9 CPU, 96GB RAM, and an NVIDIA RTX 3090ti GPU.
Figure 7[b] shows small obstacles, which we place for online learing(left) as well as our evaluation(right). In our evaluation, we place the obstacles, which is not seen in online training.
Figure 7: Overview of the prototype robot(left) [74] and small obstacles(right).
Appendix B Evaluation setup
@@
Sampling-based motion planning: This baseline generates fifteen motion primitives [75, 76] at every time step and selects the best one considering goal reaching, static and dynamic obstacles such as the pedestrians. To control the robot, we give the velocity commands corresponding to the selected motion primitive. The details are shown in the Appendix G.
@@
Residual RL [14]: In residual RL, the policy is given by the sum of a base policy and a learned policy, $a=\pi^{\text{base}}(s)+\pi^{\text{RL}}(s)$, where $\pi^{\text{base}}$ is the pre-trained control policy and $\pi^{\text{RL}}$ is the actor trained with online RL (TD3). We evaluate two choices for $\pi^{\text{base}}$: (1) the pre-trained control policy maximizing only $\sum_{t=0}^{H-1}\hat{r}^{\tt{pose}}_{t}$ to simply move towards the goal position, and (2) the pre-trained SACSoN policy, maximizing the total objective $J_{\text{sacson}}$. We label the latter as Residual RL. Residual RL provides an alternative way to combine prior policies with online model-free RL, and therefore represents a natural prior method for comparing with SELFI.
@@
All learning-based baselines use the same network structure, except that single-step methods (all except Ours and SACSoN with fine-tuning) predict only a single action $a$ rather than a sequence $\tau$. Unless specified, all RL-based methods use the same reward.
@@
In these three environments, we design the looped trajectories, which the last node is same pose as the initial node, as shown by red lines in Fig. 8, and feed the first goal image when arriving at the last node to continuously train the control policy online. We conduct online learning while rotating these loops and stop training when the robot run 15 laps or two hours. During online training, we randomly place small objects shown in Fig. 7[b](left) to learn the collision avoidance behavior for the small objects. Since Environment 1 includes many challenges for our task, we evaluate all methods in this environment and conduct the overall evaluations with the selected baselines in three environments.
Figure 8: Three environments on online training and evaluation. We conduct online training in three different challenging environments, [a] the open space facing restrooms, elevator hall and café space, [b] the entrance hall with many pedestrians, and, [c] the office area with narrow corridors. [a] and [c] have many glass walls, which are difficult for collision avoidance and cause lighting condition changes.
Appendix C Evaluation with human participants
For Q1, we conduct two types of experiments, 1) comparative evaluations with relevant baselines and 2) human subject experiments to evaluate how well our methods behave. For 1), we conduct the experiments in the “organized” environments for reproducible comparisons. For 2), we conduct our experiments in the “unorganized” environments to obtain various evaluations of each human subject’s senses. We explain the details of the experimental setups in each experiment.
C.1 Reproducible comparisons
@@
It is difficult to conduct extensive and reproducible comparisons in highly populated natural environments because pedestrians are walking in and out of the scene at random intervals. Therefore, to conduct a more controlled and reproducible comparison with each of the baselines, we focus on the “organized” environments.
Before the experiments, we instruct the human participants to follow pre-defined trajectories (as repeatably as possible) to have consistent testing conditions. However, if the robot interferes with a person’s path, the pedestrians were asked to slow down or stop as needed. If the robot can not give way to the pedestrians, the pedestrians change their path as they deem fit, to interact with the robot and then come back to the original trajectories. All subjects are asked to follow these instructions across the experiments. We conduct the experiments across different days, as well as times of day, mimicking the range of lighting conditions and changes in environment layout that a robot would experience over several days. In each experiment, we randomly place the small unseen objects shown in Appendix A into the scene to increase the clutter in the scene and evaluate the collision avoidance performance. However, we conduct the experiments under approximately equivalent conditions for each method.
C.2 Human subject experiments
To incorporate various human subjects into our evaluation, we ask the human subjects to interact with the robot and evaluate its behaviors without specifying specific ways for the subjects to interact with the robot. This prevents biased behavior and insights from the human subjects. Additionally, we ask the human subjects to have similar interactions with the robot across all methods to maintain fairness in the evaluation. To assure the behaviors of the pedestrians were as natural as possible, we did not always observe and follow the robot during online training. Subjects were not told which method was being used for each trial. Before the evaluation, we explain the rough robot route from the start to the goal position.
Appendix D Weighting objectives on online training
Similar to the other learning algorithms, our approach needs to find the best balance of each objective on online training. In our case, we first perform online learning with small weights for the learned objective $\bar{Q}(s,a)$ and gradually increase the weights. The small weighting for the learned objective makes online learning more stable and facilitates the analysis of learning results. Once we find a good weight in one environment, we use the same value in the different environments.
Appendix E Pre-training with data from the target environment
When robots operate continuously in the same environment, they can directly collect and utilize a large amount of in-domain data. In the supplemental evaluation, we show that a large offline dataset of interactions from the target environment can boost the performance of our method during online training. Specifically, we use offline RL to train our actor and critic on the collected dataset from the target environment on top of SACSoN. Then, we fine-tune both networks online by SELFI.
During offline RL training, we leverage the collected dataset from our study. While finding good hyperparameter for online learning, we collected a dataset of 50 hours of experience in Environment 1. Figure 9 shows the number of intervention for each loop in Environment 1. Even at beginning, our method leverages a large dataset, Ours + prior data only needs one intervention and can navigate the robot without any interventions after 5 laps, because of faster training with the pre-trained actor and critic on the large dataset. Similar to Table 1, Table 2 shows the mean value of the selected metrics on 5 laps in the Environment 1. Ours + prior data shows a remarkable gap against Ours in every metric except SPL. Since Ours + prior data takes a larger deviation from the original path to avoid violating the intimate distance of close pedestrians (to decrease IDV), SPL is slightly worsen.